OpenStudy (anonymous):

find the exact value of tan 2 theta given sin theta=3/5,0

OpenStudy (anonymous):

First : calculate : \[\cos\theta\] by using : \[1=\cos^2\theta+\sin^2\theta.\] Don't forget that : \[0<\theta<\frac\pi2\] so : \[\cos\theta \geq 0.\] And then : \[\tan\theta=\frac{\sin\theta }{\cos \theta}=\cdots\] And in the end use the \[\tan2\theta\] formula !